THERE ARE 850 DOUGLAS FIR AND PONDERSOA PINE TREES IN A SECTION OF FOREST BOUGHT BY LOGGING COMPANY. THE COMPANY PAID A AVERAGE OF $300 FOR EACH DOUGLAS FIR AND $225 FOR EACH PONDERSOA PINE.
IF THE COMPANY PAID $217,500 FOR THE TREES, HOW MANY OF EACH KIND DID THE COMPANY BUY?
THANKS
JMS

Answer by checkley71(8403)

300X+225(850-X)=217,500
300X+191,250-225X=217,500
75X=217,500-191,250
75X=26250
X=26,250/75
X=350 DOUGLAS FIR TREES.
850-350=500 PONDEROSA PINES.
PROOF
300*350+225*500=217,500
105,000+112,500=217,500
217,500=217,500